I know :
GDP 18000
T 4000
C 7600
I 4500
G 5000
NX 4500
==================
How can i find the equilibrium level of income ?
How can i find UI value ?

Thanks.
1
Expert's answer
2014-05-06T08:38:07-0400
Equilibrium income refers to the state at whichaggregate quantity supplied is equal to aggregate quantity demanded. It is
usually stable provided that the various factors involved do not change. It is calculated using the formula, GDP = C + I + G + (X - M), where C is consumption, I is investments, G is government spending and X-M is exports
minus imports. This can be formulated when one knows the parameters to be used in the calculation.

So, GDP = 7600 + 4500 + 5000 + 4500 = 21600 is equilibrium level of income.